# Understanding Common Ground Misalignment in Goal-Oriented Dialog: A Case-Study with Ubuntu Chat Logs

### v1

@@ 7 Related Work
Markowska et al. (2023) track speaker versions of the CG through speaker “beliefs” expressed in conversations in the LDC Callhome Canavan et al. (1997) corpus. However, participants are not as incentivized to build and maintain a rich CG since the conversations are not goal-driven, and are between close friends or family. Khebour et al. (2024) annotate a task-oriented corpus for multi-modal features and dialogue moves to model shared beliefs and questions under discussion. The authors train LSTM-based classifiers of dialogue moves relevant to tracking CG, finding that utterances may or may not be aligned with other modalities such as gesture. This highlights the challenge of tracking common ground in physically situated dialogue; our dataset simplifies focus to text alone.
Shaikh et al. (2024) use grounding acts to compare the degree of grounding by LLMs in human-LLM conversations, finding that LLMs perform less computational grounding. Our work complements these directions by focusing on whether LLMs can detect when and how participants in a conversation might lose track of CG.
@@ A.2 Model Finetuning Details
For our finetuning experiments, we create five random train-test splits of our dataset, with 30% of conversations in Ubuntu-CG in each split going towards the “test” set. For each of our context windows ($k=3$ or $k=5$ turns), we train a distilroberta-base model with 82M parameters for 15 epochs with a learning rate of $54e-$ 5. As evaluation data, we pick a single fold and isolate a part of its training set as our development data. Since this is a class-imbalanced dataset (turns not containing friction greatly outnumber turns containing friction 10 to 1), we artificially reduce the number of negative samples in our training data, however resampling was not done on the test data. We also experimented with larger encoder-only models, but they all performed much poorer than distilroberta under both full-parameter and classification head-only training.
@@ A.4 Prompts
A strong indicator of conversational friction could be a participant asking the other participant to revisit or clarify previously shared content in the conversation, in a process known as conversational repair. However, in many cases there may not be an explicit Repair Request issued by a participant but from the context it can be reasoned that a participant is struggling to keep up with the conversation.
NOTE: Friction is often signaled by the helpee asking a followup question. However, not all followup questions indicate that the speakers are not on the same page. For example, clarification questions that ask for information not assumed by either user to be in the common ground are not cases of conversational friction. Clarification questions that move the conversation forward without questioning the common ground are not cases of conversational friction. If there is **no conversational friction** make sure to indicate that in the output by setting "friction_present" to false.
@@ A.6 Annotator Instructions
Before any annotation task, annotators had to fill-up a consent form (Figure 3). To ensure we’re measuring equivalent constructs, the annotator instructions was kept identical to Prompt A.4. A more detailed instruction document can be found in the supplementary material. The similarity scoring prompt is shown in Figure 4.
@@
A.7 Metalinguistic vs Linguistic Capabilities of LLMs
LLMs perform differently in a linguistic task (responding to a query) as opposed to a metalinguistic task (spotting a mismatch in common ground) between participants. It is possible that a model response addresses a mismatch in common ground indirectly as a conversational partner while failing to identify the cause of friction as an observer. However, we believe that understanding LLM behavior of tracking common ground is an essential prerequisite to many other downstream research questions, such as cases where an LLM is used as a conversational facilitator Argyle et al. (2023).
A.8 Breakdown of Table 5

### v2

@@ 7 Related Work
Markowska et al. (2023) track speaker versions of the CG through speaker “beliefs” expressed in conversations in the LDC Callhome Canavan et al. (1997) corpus. However, participants are not as incentivized to build and maintain a rich CG since the conversations are not goal-driven, and are between close friends or family. Khebour et al. (2024) annotate a task-oriented corpus for multi-modal features and dialogue moves to model shared beliefs and questions under discussion. The authors train LSTM-based classifiers of dialogue moves relevant to tracking CG, finding that utterances may or may not be aligned with other modalities such as gesture. This highlights the challenge of tracking common ground in physically situated dialogue; our dataset simplifies the focus to text alone.
Shaikh et al. (2024) use grounding acts to compare the degree of grounding by LLMs in human-LLM conversations, finding that LLMs perform less computational grounding. Our work complements these directions by focusing on whether LLMs can detect when and how participants in a conversation might lose track of CG. In more recent work,  Shaikh et al. (2025) focuses on this divergence across several grounding acts, and introduces an annotated dataset of LLMs failing to ground in human-LLM conversations.
Also related to our work is the concept of positive friction as explored in  İnan et al. (2025). Here, the authors view “friction” not as a misalignment in common ground that detracts from the goal of the conversation, but as a series of communicative “movements” such as pausing, revealing speaker assumptions, etc., that facilitate long-term success over short-term progress in a conversation.
@@ 8 Conclusion and Future Work
Another limitation of our work stems from the fact that we limit our analyses of LLMs to their role as an observer of human-human conversation, and not as a participant. Given that LLMs perform differently in linguistic tasks (such as responding to a query) as opposed to metalinguistic tasks (spotting a mismatch in common ground), it is possible that a model response addresses a mismatch in common ground indirectly as a conversational partner while failing to identify the cause of friction as an observer Hu and Frank (2024). However, we believe that understanding LLM behavior of tracking common ground is an essential prerequisite to many other downstream research questions, such as cases where an LLM is used as a conversational facilitator Argyle et al. (2023).
@@
Acknowledgments
We thank our anonymous reviewers for their comments and suggestions throughout the reviewing process. We also thank Atrey Desai, Nyle Masood, Ayush Kumar, Maya Srikanth, and Pratheek Hegde for providing valuable annotations. This work was supported in part by the US National Science Foundation award 2124270. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation.
@@ A.2 Model Finetuning Details
[…]
@@ A.4 Prompts
A strong indicator of conversational friction could be a participant asking the other participant to revisit or clarify previously shared content in the conversation, in a process known as conversational repair. However, in many cases, there may not be an explicit Repair Request issued by a participant, but from the context, it can be reasoned that a participant is struggling to keep up with the conversation.
NOTE: Friction is often signaled by the helpee asking a follow-up question. However, not all follow-up questions indicate that the speakers are not on the same page. For example, clarification questions that ask for information not assumed by either user to be in the common ground are not cases of conversational friction. Clarification questions that move the conversation forward without questioning the common ground are not cases of conversational friction. If there is **no conversational friction** make sure to indicate that in the output by setting "friction_present" to false.
@@ A.6 Annotator Instructions
Before any annotation task, annotators had to fill out a consent form (Figure 3). To ensure we’re measuring equivalent constructs, the annotator instructions was kept identical to Prompt A.4. A more detailed instruction document can be found in the supplementary material. The similarity scoring prompt is shown in Figure 4.
@@
A.7 Breakdown of Table 5
@@
A.8 Macro-Averaged Results
For completeness, we also present the macro-averaged result in Table 11.
Friction Found Friction Overlap #Predictions
Model Precision Recall F1 Precision Recall F1
gpt-4o 34.34 71.43 46.38 14.28 29.70 19.29 495
gpt-4o w/ Elab. 34.94 63.87 45.17 14.75 26.96 19.07 435
gpt-4o-mini 32.28 42.86 36.82 13.97 18.55 15.94 316
gpt-4o-mini w/ Elab. 28.32 46.64 35.24 14.11 23.23 17.55 392
Llama-3.1-8b-Instruct 14.82 79.83 25.00 5.88 31.68 9.92 1282
Llama-3.1-8b-Instruct w/ Elab. 15.88 83.61 26.69 6.37 33.53 10.70 1253
Llama-3.1-70b-Instruct 22.40 80.67 35.07 9.06 32.63 14.18 857
Llama-3.1-70b-Instruct w/ Elab. 16.89 68.07 27.07 7.08 28.54 11.35 959
distilroberta-base ($k=3$, finetuned) 19.89 48.07 27.99 7.70 18.36 10.80 -
distilroberta-base ($k=5$, finetuned) 18.32 46.43 26.16 8.42 20.97 11.97 -
Table 11: Macro-averaged precision, recall, and F1 scores of different models on detecting friction. Each friction instance contributes equally to the final metrics regardless of which conversation it appears in. #Predictions refer to the total number of instances of conversational friction found by each model. For reference, annotators identified 238 instances in total. gpt-4o without Elaboration of technical terms (Sec 3.1) performed best across all models. Micro-averaged results are provided in the main text (Table 6).